OpenStudy (anonymous):

Solve the absolute value equation. 2 | 3x-5 | - 8 = 8

OpenStudy (anonymous):

the numbers inside the | | mean absolute value

OpenStudy (anonymous):

which means keep it positive

OpenStudy (anonymous):

soo its gonna look like 2(3x+5)-8=8

OpenStudy (anonymous):

the answer turns into 6x+5-8=8 then x=5/6

OpenStudy (anonymous):

thanks

OpenStudy (yttrium):

No. you should have two values.

OpenStudy (anonymous):

is it x= -1 5/6

OpenStudy (yttrium):

Considering the negative and positive values inside the abs value. we have \[2|3x-5| - 8 = 8\] upon simplifying\[|3x-5| = 8\] then split it into two cases 3x-5 = 8 -(3x-5) = 8 Then solve for x, there you can find the answer.

OpenStudy (yttrium):

The answer should be x = 13/3 and -1

OpenStudy (anonymous):

so whats the right answer?

OpenStudy (yttrium):

x = 13/3 and x=-1